OpenStudy (anonymous):

work out values of a & b if x^2+ax-2=(x-3)^2-b

OpenStudy (anonymous):

a=-6 b=11

jhonyy9 (jhonyy9):

x2 +ax -2=x2 -6x +9 -b x2 +ax +b=x2 -6x +9+2 x2 +ax +b=x2 -6x +11